OpenStudy (anonymous):

(a) Factorise x^2-9. (b) Use your answer to part (a) to find two factors of 9991, other than 1 and 9991.

OpenStudy (anonymous):

I know (A) is (x-3)(x+3). But I don't know how to do part (b).

OpenStudy (anonymous):

9991=(x+3)(x-3)=x^2-9 10000=x^2 x=100 9991=(100+3)(100-3) 9991=103*97

OpenStudy (experimentx):

Oh ... lol sorry

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Sure thing :)